Siphesihle borrows money from the bank at a dis ount rate of 16,5% per year. He must pay the bank R30 000 in eight months from now. The amount of money he re eives from the bank now is

[1]R33 707,87

. [2} R26 700,00.

[3} R27 027,03.

[4} R33 463,26.

[5} R33 300,00. 


1
Expert's answer
2021-06-07T17:27:08-0400

Time, t = 8 months = 8/12 year = 2/3 year

Discount rate, r = 16.5% per year = 0.165

Let M be the borrowed money from the bank.

And D be the discounted amount.

Then, D=M.r.t

D=M(0.165)(2/3)

D=0.11M ...(A)

Now, P=30000

P=M-D

30000=M-0.11M [Using (A)]

0.89M=30000

M=33,707.87

Thus, option 1, i.e., R33707.87 is correct